solve the following equations for x and y by elimination
3x + 3y=0
2x - 2y=8

Answer by jennywecas(28)

2*(3x + 3y=0)=6x+6y=0
3*(2x - 2y=8)=6x-6y=24
12x=24
x=2
3(2)+3y=0
6+3y=0
3y=-6
y=-2
Answer is (2,-2)